MakaylaChuck23:

Solve for 𝑥 x: 3 𝑥 + 5 = 2 𝑥 + 9 3x+5=2x+9

MakaylaChuck23:

can someone explain?

SaltySuga:

1. Start with the equation 3x + 5 = 2x + 9 2. Get all the x terms on one side To do this, you can subtract 2x from both sides: 3x - 2x + 5 = 9 This simplifies to: x + 5 = 9 3. Isolate x Now, subtract 5 from both sides to solve for x: x = 9 - 5 Which gives you: x = 4 The solution is x = 4

MakaylaChuck23:

Thank you.

NFangStar:

Awesome Job @makaylachuck23